HIGH RESOLUTION INFRARED AND MICROWAVE STUDY OF ^{10}BF_{2}OH AND ^{11}BF_{2}OH: the 5^{1}, 6^{1}, 7^{1}, 8^{1}, 9^{1} AND 8^{1}9^{1} VIBRATIONNALLY EXCITED STATES.

\underline{A. PERRIN}, M. CARVAJAL-ZAERA, Z. DUTKIEWICZ, Laboratoire de Photophysique Moléculaire, CNRS, Université Paris Sud, Campus d’Orsay, Bat 350, 91405 Orsay Cedex, France; J.- M.FLAUD, Laboratoire Interuniversitaire des Systemes Atmosphériques (LISA), CNRS, Université Paris 12, Bat. P1 61, avenue du Général de Gaulle 94010 Créteil Cedex, France; H. BÜRGER, D.COLLET, Anorganische Chemie, FB 9, Universität-Gesamthochschule, D-42097 Wuppertal, Germany; J.DEMAISON, F.WILLAERT, Laboratoire PhLAM, CNRS, Université de Lille I, Bat. P5, 59655 Villeneuve d’Ascq Cedex, France; H. MÄDER, Institut für Physikalische Chemie, Olshausenstrasse 40, D-24098 Kiel, Germany; N. W. LARSEN, Chemical Laboratory 5, H.C. Oersted Institute, 2100 Copenhagen, Denmark.

High resolution ( 2-3 x 10^{−3} cm^{−1}) Fourier transform infrared spectra of gas phase ^{10}B and ^{11}B enriched isotopic and natural samples of BF_{2}OH (difluoroboric acid) were recorded in the 400- 4000 cm^{−1} spectral range. Starting from the results of a previous study^{a} which involved the ν_{8} (BF_{2} out-of-plane bending) and ν_{9} (OH torsion) bands of ^{11}BF_{2}OH, it has been possible to perform the first rovibrational analysis of the ν_{5} (BF_{2} bending), ν_{8}, ν_{9} and ν_{8}+ν_{9} bands of ^{10}BF_{2}OH, and of the ν_{5}, ν_{7} (F_{2}BO in-plane bending) and ν_{8}+ν_{9} bands of ^{11}BF_{2}OH up to very high rotational quantum numbers. In addition microwave transitions within the 5^{1}, 6^{1}, 7^{1}, 8^{1} and 9^{1} vibrational states of ^{11}BF_{2}OH were measured using predictions performed from ab initio calculations. For the ν_{5} and ν_{7} bands of ^{11}BF_{2}OH, C-type Coriolis resonances coupling the 5^{1} and 7^{1} energy levels with those of the 7^{2} and 6^{1} dark states respectively were accounted for in the calculations. The other bands appeared to be unperturbed,

aD.Collet, A.Perrin, H.Bürger, and J.-M.Flaud, J. Mol. Spectrosc. 212, 118 (2002)